OpenStudy (anonymous):

BC calc question: Find the area of the regions enclosed by the lines and curves y=x(sqrt(a^2-x^2)), where a>0, and y=0

OpenStudy (anonymous):

I have tried this problem so many times and so many ways that I came here. I tried integration by parts, u-sub, etc. Any help is appreciated

OpenStudy (anonymous):

Let me clarify, I know what to do, but it comes down to integrating the first function one way or the other

OpenStudy (anonymous):

And I have no idea how to integrate that function

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Well i think that it's a common integral

OpenStudy (anonymous):

Well I've tried solving it. Can you walk through it?

OpenStudy (anonymous):

\[\int\limits \sqrt{a^2-x^2}dx\] right?

OpenStudy (anonymous):

x times that

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

even easier

OpenStudy (anonymous):

u sub u=x^2, have you tried that?

OpenStudy (anonymous):

I didn't know how to do it with the two variables

OpenStudy (anonymous):

a is a constant

OpenStudy (anonymous):

WHAT

OpenStudy (anonymous):

I spent SO LONG on this problem trying to figure out how to do this

OpenStudy (anonymous):

yup you don't learn double integration in bc iirc

OpenStudy (anonymous):

no a is definitely a constant

OpenStudy (anonymous):

Hahah thanks man

OpenStudy (anonymous):

np, good luck :)

OpenStudy (anonymous):

Ok I'm still running into problems. Can you run through it with me really quickly?

OpenStudy (anonymous):

Wait actually if a is a constant, then when you derive, it goes to 0. Nevermind!